# Enhancing LLMs for Power System Simulations: A Feedback-driven Multi-agent Framework

### v2

@@ V-E Cost Analysis
It is noteworthy that aside from parallel retrieval, no specialized acceleration techniques were employed in this framework. Thus, despite its already satisfactory performance, there is considerable potential for speed enhancements. Even in the current state, without any specific acceleration strategies, GPT4o-Full can execute approximately 120 simulation tasks per hour in Daline and MATPOWER, with a total token cost of around 1.68 USD. Given this high efficiency and cost-effectiveness, the proposed framework presents a promising pathway to improve researcher productivity.
@@
VI Conclusion
This paper addresses the research gap in enhancing LLMs for power system simulations by proposing a feedback-driven, multi-agent framework. It represents the first systematic approach to significantly improve LLMs’ simulation capabilities across both familiar and new tools. Validated on 69 diverse simulation tasks from Daline and MATPOWER, the proposed framework achieved substantial performance improvements, with success rates of 93.13% and 96.85%, respectively. It far surpasses those of baseline schemes, including the latest LLM, o1-preview, and SFT for LLMs. Key findings include: (i) The original simulation capability of LLMs is limited, as evidenced by GPT4o and o1-preview achieving success rates no higher than 27.77%. (ii) Even with the standard RAG module and a comprehensive knowledge base, LLMs achieve overall success rates below 45%, highlighting the need for a more comprehensive approach. (iii) Our framework’s high success rate stems from a synergistic integration of enhanced RAG, enhanced reasoning, as well as environmental acting and feedback mechanisms. Removing any of these elements results in a significant performance decline. (iv) While SFT may be able to capture coarse-grained attributes such as style, tone, and format, it struggles to retain the full detail necessary for power system simulations. Even when the entire test dataset is included in the training set, SFT achieves a success rate of less than 60%. This limitation stems from its inherently lossy compression mechanism and constrained parameter update capacity, which prevent it from encoding fine-grained details with high fidelity. (iv) The proposed framework enables LLMs to execute tasks efficiently, with each task completed in approximately 30 seconds at a token cost of only 0.014 USD (for the simulation tasks used in this paper), offering a scalable, cost-effective solution that enhances productivity of human scientists in power systems.
@@ VI Conclusion
However, several critical future challenges still remain. First, while the proposed framework demonstrates significant improvements over existing schemes and supervised fine-tuning, achieving a perfect 100% success rate for LLM-based simulations remains an open challenge. For instance, the proposed framework sometimes fails to detect “non-execution-bug” failures, resulting in unaware inaccurate outputs that cannot be automatically corrected. A potential avenue for future work is to explore the use of a RAG-supported, fine-tuned LLM-based code-checking agent, which would verify executed simulation code and identify hidden errors that the error-reporting mechanism may overlook. Second, when simulation requests are inherently ambiguous or underspecified, accurately determining the user’s intent becomes difficult for any automated method—not just the framework presented in this paper. This ambiguity primarily arises from the initial lack of detail in the user’s input, rather than any intrinsic limitation of natural language coding. To address this, a promising future direction is to integrate an interactive dialogue stage into the framework. In this approach, an autonomous agent would evaluate the specificity of the initial request and, if necessary, prompt the user with targeted clarifying questions. This human-in-the-loop strategy could help refine the requirements, leading to more precise retrieval and improved code generation.

### v3

@@ V-F Cost Analysis
It is noteworthy that, aside from parallel retrieval, no specialized acceleration techniques were employed in this framework. Thus, despite its already satisfactory performance, there is considerable potential for speed enhancements. Even in the current state, without any specific acceleration strategies, GPT4o-Full can execute approximately 120 simulation tasks per hour in Daline and MATPOWER, with a total token cost of around 1.68 USD. Given this high efficiency and cost-effectiveness, the proposed framework presents a promising pathway to improve researcher productivity.
@@
VI Conclusion
This paper addresses the research gap in enhancing LLMs for power system simulations by proposing a feedback-driven, multi-agent framework. It represents the first systematic approach to significantly improve LLMs’ simulation capabilities across both familiar and new tools. Validated on 69 diverse simulation tasks from Daline and MATPOWER, the proposed framework achieved substantial performance improvements, with success rates of 93.13% and 96.85%, respectively. It far surpasses those of baseline schemes, including the latest LLM, o1-preview, and SFT for LLMs. Key findings include: (i) The original simulation capability of LLMs is limited, as evidenced by GPT4o and o1-preview achieving success rates no higher than 27.77%. (ii) Even with the standard RAG module and a comprehensive knowledge base, LLMs achieve overall success rates below 45%, highlighting the need for a more comprehensive approach. (iii) Our framework’s high success rate stems from a synergistic integration of enhanced RAG, enhanced reasoning, as well as environmental acting and feedback mechanisms. Removing any of these elements results in a significant performance decline. (iv) While SFT may be able to capture coarse-grained attributes such as style, tone, and format, it struggles to retain the full detail necessary for power system simulations. Even when the entire test dataset is included in the training set, SFT achieves a success rate of less than 60%. This limitation stems from its inherently lossy compression mechanism and constrained parameter update capacity, which prevent it from encoding fine-grained details with high fidelity. (iv) The proposed framework enables LLMs to execute tasks efficiently, with each task completed in approximately 30 seconds at a token cost of only 0.014 USD (for the simulation tasks used in this paper), offering a scalable, cost-effective solution that enhances the productivity of human scientists in power systems.
Overall, the effectiveness of this work relies on the coordinated integration of general-purpose LLM capabilities (such as natural language understanding and code generation) with carefully designed domain-specific strategies tailored to power system simulations. These strategies are essential for addressing the unique challenges posed by simulation tasks. While this work focuses on simulations, the proposed framework is inherently designed to enable LLMs to adapt to unfamiliar domains and is readily extensible to broader power system research. Furthermore, the proposed simulation-enhanced LLM system may serve as a key component in future human-machine collaborative research, supporting not only simulation execution but also the validation and refinement of research ideas.
However, several critical future challenges still remain. First, while the proposed framework demonstrates significant improvements over existing schemes and supervised fine-tuning, achieving a perfect 100% success rate for LLM-based simulations remains an open challenge. For instance, the proposed framework sometimes fails to detect “non-execution-bug” failures, resulting in unaware, inaccurate outputs that cannot be automatically corrected. A potential avenue for future work is to explore the use of a RAG-supported, fine-tuned LLM-based code-checking agent, which would verify executed simulation code and identify hidden errors that the error-reporting mechanism may overlook. Second, when simulation requests are inherently ambiguous or underspecified, accurately determining the user’s intent becomes difficult for any automated method, not just the framework presented in this paper. This ambiguity primarily arises from the initial lack of detail in the user’s input, rather than any intrinsic limitation of natural language coding. To address this, a promising future direction is to integrate an interactive dialogue stage into the framework. In this approach, an autonomous agent would evaluate the specificity of the initial request and, if necessary, prompt the user with targeted clarifying questions. This human-in-the-loop strategy could help refine the requirements, leading to more precise retrieval and improved code generation.